# Classification of $f(R)$ Theories Of Inflation

### v1

@@ 1 Introduction
Inflation is one of the best paradigm used to solve the main puzzles of primordial cosmology: the horizon problem and the flatness problem. In particular, it strikes for its simplicity. Indeed, it is sufficient to extend Einstein gravity with the $R^{2}$ term to explain the present data and solve the above problems, as shown by Starobinsky [1]. Alternatively, the same result can be achieved by introducing an additional scalar field with a suitable potential [2, 3], without modifying general relativity.
@@
During the last decades these two possibilities has been largely studied and extended by considering different potentials [4] or by generalizing Starobinsky model to the so-called $f(R)$ theories [5, 6], where a general function of the scalar curvature $R$ is used in the gravitational action. Furthermore, the same models have been further extended by using the Palatini formulation of gravity, that for a certain class of theories, gives different physical predictions compared to the standard metric formulation (see e.g. [7] and [8] for a review). In the end, the simplicity of the solution offered by inflation has been traded for a plethora of models that are all compatible with the data and suitable to solve the open problems. Therefore, until we do not have other strong data to falsify most of these models, it is compulsory to include some fundamental principles that can help us navigate and select them out. A good lesson can be learned from particle physics, where power counting, renormalizability, unitarity and gauge symmetries strongly constrain the models. In this regards, it is remarkable that Starobinsky inflation is obtained by including the only term with dimension 4 that can be added to the Hilbert action without spoiling unitarity. This points in the direction that power counting, renormalizability and unitarity might play important roles in primordial cosmology. It is also possible to make Starobinsky model renormalizable by including in the gravitational action all terms with dimension 4 [9]. This includes the cosmological term and the $R_{\mu\nu}R^{\mu\nu}$ term. However, the latter causes a violation of unitarity due to the presence of a massive spin-2 ghost. Such unwanted degree of freedom can be removed freedom the spectrum without renouncing to renormalizability by turning it into a so-called “purely virtual” particle [10, 11], thus restoring unitarity. However, in the context of cosmology the presence of the $R_{\mu\nu}R^{\mu\nu}$ term affects only the perturbations [12], while the background dynamics stays unmodified. Nevertheless, a sharp prediction for the tensor-to-scalar ratio has been derived [12].
@@
Recently, a new way of classify single-field inflationary models has been proposed [13]. It is based on a mathematical analogy between the renormalization-group (RG) flow in quantum field theory and the evolution of the background metric during inflation [14], where the role of the coupling constant is played by an appropriate slow-roll parameter $\alpha$. In this framework, a key quantity is the analogous of the beta function $\beta_{\alpha}$, which encodes the background dynamics in a perturbative way. In particular, it is required to be a power series in $\alpha$ that starts from a quadratic term with negative coefficient. Then, the various scalar-field potentials can be classified based on their beta functions. However, the study of $f(R)$ was left out. The reason is that, moving to the Einstein frame, each $f(R)$ generates a different and rather complicated potential and, in some cases, it is not even possible to determine them analytically. Moreover, since the definition of $\alpha$ is arbitrary, in the Einstein frame it is nontrivial to find the right one to study the $f(R)$ altogether. The only exception is the potential associated to Starobinsky model, which was indeed included in the analysis of [13].
@@
In this paper we extend the classification of [13] to $f(R)$ theories directly in the Jordan frame, which allows us to identify the correct definition for $\alpha$. In particular, assuming that the beta function be a power series, we classify the models and derive the leading predictions class by class without specifying the form of $f(R)$. Moreover, starting from a beta function with generic coefficients, we work out the expression of all the $f(R)$ that belong to the same class in terms of such coefficients.
@@
Comparing the results with the most up-to-date experimental constraints we find conditions on the first nonvanishing coefficient of the beta functions and exclude all classes but two. Remarkably, if we focus on polynomial $f(R)$, which in general relativity are generated by renormalization anyway, we find that the only one suitable for slow-roll inflation is Starobinsky model. In fact, extending the analysis to theories with a linear beta function, it turns out that all the other polynomials belong to this category. A particular property of these theories is that they cannot realize slow-roll inflation but only constant-roll inflation [15], i.e. an accelerated expansion with a constant rate. We highlight some difficulties in interpolating between slow-roll and constant-roll inflation, and, thus, between Starobinsky model and its polynomial extensions.
@@
Finally, once we derived all the quantities in the Jordan frame we provide the map to the Einstein frame for comparison with [13].
@@
The paper is organized as follows. In section 2 we review $f(R)$ theories, introduce the RG analogy and find the definition of $\alpha$ that simplifies the equations. In section 3 we derive the quadratic action for tensor and scalar perturbations in a generic $f(R)$ theory and recall the derivation of the spectra used in [13]. In section 4 we work out the leading-order predictions for the spectral indices and the tensor-to-scalar ratio and show that every class that does not have quadratic or cubic beta function is ruled out by data. In section 5 we derive the general form of the $f(R)$ that belong to the surviving classes and their power spectra. In section 6 we derive the perturbative map between the Jordan frame and the Einstein frame. In section 7 we study a special class of theories with linear beta function and point out the issues with these models. Finally, section 8 contains our conclusions.
@@
Notation and conventions: We use the signature (+,,,)(+,-,-,-)( + , - , - , - ) for the metric tensor. The Riemann and Ricci tensors are defined as $R^{\mu}_{\ \nu\rho\sigma}=\partial_{\rho}\Gamma^{\mu}_{\nu\sigma}-\partial_{\sigma}\Gamma^{\mu}_{\nu\rho}+\Gamma^{\mu}_{\alpha\rho}\Gamma^{\alpha}_{\nu\sigma}-\Gamma^{\mu}_{\alpha\sigma}\Gamma^{\alpha}_{\nu\rho}$ and $R_{\mu\nu}=R^{\rho}_{\ \mu\rho\nu}$, respectively. In the case of cosmological perturbations we omit the integral over the space momentum and use the word “action” for the integral over time of the lagrangian density.
@@ 2 $f(R)$ theories and the cosmic RG flow
We start by recalling the action and equations of motion on Friedmann-Lemaître-Robertson-Walker (FLRW) background. The starting action reads
@@
where $G$ is Newton constant. The equations of motion are
@@
where $H\equiv\dot{a}/a$ is the Hubble constant. Now we adapt the procedure of [13] to $f(R)$ theories. The ingredients to define a RG flow are
@@
    a first order differential equation for that small quantity (analogous of the beta function);
@@
Indeed, this is a good choice if we work with a specific $f(R)$ model (see [12]). However, we would like to treat these type of theories generically. It turns out that another useful quantity is instead
@@
which we show below to be small during inflation. Following [13] we also define
@@
$-v\beta_{\alpha}=\alpha(1+2\alpha)-(1-\alpha^{2})\varepsilon,$ (2.10)
@@
where the $b_{i}$ are constant coefficients. Note that if we admit a linear beta function, then $\varepsilon=\mathcal{O}(\alpha)$, while $\eta_{H}=\mathcal{O}(\alpha^{0})$. In this case slow-roll inflation cannot be realized because $\eta_{H}$ is constant in the far past. This class of models are discussed in section 7.
As mentioned above, we could have chosen $\varepsilon$ as small quantity to define the RG flow. In that case we would have had
@@
Again, assuming $\varepsilon\ll 1$, we need that $\beta_{\varepsilon}$ be at least quadratic to guarantee $\eta_{H}=\mathcal{O}(\varepsilon)$. Since $\varepsilon(\alpha)$ is a simple power series, it means that also $\alpha(\varepsilon)$ is. We conclude that assuming $\alpha\ll 1$ and $\beta_{\alpha}=\mathcal{O}(\alpha^{2})$ is equivalent to assume $\varepsilon\ll 1$ and $\beta_{\varepsilon}=\mathcal{O}(\varepsilon^{2})$. This means that, during the inflationary phase, the roles of $\alpha$ and $\varepsilon$ are interchangeable without loss of generality. Passing from using one or the other is just a perturbative redefinition.
Finally, the background spacetime needs to reach the de Sitter one in the asymptotic past. This means that the first nonvanishing coefficient of the beta function must be negative. Because of this properties, the whole construction resembles that of beta functions and running-coupling constants in quantum field theory, where an asymptotically-free theory requires a negative beta function. Here the “flow” is represented by the background evolution during the perturbative phase of inflation. More details can be found in [13]. It is important to highlight that such an analogy is just a mathematical correspondence and has nothing to do with quantum corrections, which are not considered in this paper.
@@
The equation (2.17) can be solved by a simple power series
@@
The equation (2.25) can be solved by truncating the beta function. Since its formal expressions do not depend on the frame, we can use the same listed in [13]. For class I and II at the leading-log order (LLO) we have
@@
We do not report the next-to-leading-log (NLL) corrections since in this paper we can directly use the results of [13]. However, it is important to highlight that all the results for power spectra and spectral indices will eventually be written as expansions in $\alpha_{k}$, which is fixed at a pivot scale $k_{*}$ by the experimental data. Then thanks to the second equation in (2.26) we can evolve $\alpha$ to an arbitrary $k$. For example, in class I we have
@@
Since $\mathcal{P}$ can be viewed as a function of $\alpha(\ell)$ and $\alpha_{k}$, its conservation on superhorizon scales means that the dependence on $\alpha$ drops and $k$ appears only through $\alpha_{k}$. More detail can be found in [14].
@@ 3 Perturbations
In this section we outline the general strategy to derive the quadratic action for the perturbations and solve their equations of motion by briefly reviewing the method used in [13]. However, we can directly export the solutions of the Mukhanov-Sasaki equation as well as the expressions for the power spectra and spectra indices at the NNLL and NLL for tensor and scalar perturbations, respectively. The reason is that in [13] those results were derived for a general action quadratic in the a perturbation, without specifying its origin.
@@
where $\sigma$ is a function that depends on background quantities. Typically [13], $\sigma$ is analytic in $b_{i}$, which means that it can only depends on $v$ and $\beta_{\alpha}$, as well as on $\alpha$ explicitly, because $H$, $F$, $f$ and $R$ are different for each class and not analytic in the first nonvanishing $b_{i}$.
@@
where $g_{n}$ are functions that depends on $\sigma$ and on the $w_{m}$ with $m<n$. The Bunch-Davies condition
@@
Finally, the solution in terms of $f_{n}$ is [14]
@@
The procedure just explained gives the $w$ as an explicit function of $\eta$. However, it is possible to single out the part of $w$ (more precisely of $\eta w$) that survives the superhorizon limit. Such part depends on $\ln\eta$ only. For this reason, it can be written as a function of $\alpha$. We summarize the steps here below. More details can be found in [14] and  [13].
@@
and the equation (3.9) turns into
@@
Both in the tensor and scalar case, the constant $C$ can be chosen such that $w^{\text{C}}$ coincides with the usual gauge-invariant perturbations, which are also conserved in the superhorizon limit. The power of this method is that it gives a recipe to obtain gauge-invariant perturbations that are conserved on superhorizon scales by construction. Although this result might not seems surprising in the case of single-field inflation, it becomes particularly relevant in the case of multifield inflation, where it is nontrivial to find conserved perturbations. See [16] and [17] for applications of the RG techniques in the context of double-field inflation.
@@
As anticipated in the beginning of this section, since in [13] the functions $w_{n}$, as well as $J$ and $Q_{0}$, have been computed assuming a power series for $\sigma$ with generic coefficients, we can directly use those results. The expressions for $w_{n}$ are analytic in $b_{i}$, while those for $J$, and therefore those for power spectra, depend on the class. The reason is that the transformation from $w$ back to the original perturbation involves $F$, $H$ and $f$.

### v2

@@ 1 Introduction
Inflation is one of the best paradigms used to solve the main puzzles of primordial cosmology: the horizon problem and the flatness problem. In particular, it strikes for its simplicity. Indeed, it is sufficient to extend Einstein gravity with the $R^{2}$ term to explain the present data and solve the above problems, as shown by Starobinsky [Starobinsky:1980te]. Alternatively, the same result can be achieved by introducing an additional scalar field with a suitable potential [Guth:1980zm, Linde:1981mu], without modifying general relativity.
@@
During the last decades these two possibilities have been largely studied and extended by considering different potentials [Martin:2013tda] or by generalizing the Starobinsky model to the so-called $f(R)$ theories [Sotiriou:2008rp, DeFelice:2010aj], where a general function of the scalar curvature $R$ is used in the gravitational action. Moreover, the same models have been further extended by using the Palatini formulation of gravity that, for a certain class of theories, gives different physical predictions compared to the standard metric formulation (see e.g. [Bauer:2008zj] and [Olmo:2011uz] for a review). In the end, the simplicity of the solution offered by inflation has been traded for a plethora of models that are all compatible with the data and suitable to solve the open problems. Therefore, until we do not have other strong data to falsify most of these models, it is compulsory to include some fundamental principles that can help us navigate and select theories out. A good lesson can be learned from particle physics, where power counting, renormalizability, unitarity and gauge symmetries strongly constrain the models. In this regard, it is remarkable that Starobinsky inflation is obtained by including the only term with dimension 4 that can be added to the Hilbert action without spoiling unitarity. This points in the direction that power counting, renormalizability and unitarity might play important roles in primordial cosmology. It is also possible to make the Starobinsky model renormalizable by adding the gravitational action all terms with dimension 4 [Stelle:1976gc]. This includes the cosmological term and the $R_{\mu\nu}R^{\mu\nu}$ term. However, the latter causes a violation of unitarity due to the presence of a massive spin-2 ghost. Such unwanted degree of freedom can be removed from the spectrum without renouncing to renormalizability by turning it into a so-called “purely virtual” particle [Anselmi:2018ibi, Anselmi:2018tmf], thus restoring unitarity. However, in the context of cosmology the presence of the $R_{\mu\nu}R^{\mu\nu}$ term affects only the perturbations [Anselmi:2020lpp], while the background dynamics stays unmodified. Nevertheless, a sharp prediction for the tensor-to-scalar ratio has been derived [Anselmi:2020lpp].
@@
Recently, a new way to classify single-field inflationary models has been proposed [Anselmi:2021rye]. It is based on a mathematical analogy between the renormalization-group (RG) flow in quantum field theory and the evolution of the background metric during inflation [Anselmi:2020shx], where the role of the coupling constant is played by an appropriate slow-roll parameter $\alpha$. In this framework, a key quantity is the analogous of the beta function $\beta_{\alpha}$, which encodes the background dynamics in a perturbative way. In particular, it is required to be a power series in $\alpha$ that starts from a quadratic term with negative coefficient. Then, the various scalar-field potentials can be classified based on their beta functions. However, in in [Anselmi:2021rye] the study of $f(R)$ was left out. The reason is that, moving to the Einstein frame, each $f(R)$ generates a different and rather complicated potential and, in some cases, it is not even possible to determine them analytically. Moreover, since the definition of $\alpha$ is arbitrary, in the Einstein frame it is nontrivial to find the right one to study the $f(R)$ altogether. The only exception is the potential associated to the Starobinsky model, which was indeed included in the analysis of [Anselmi:2021rye].
@@
In this paper we extend the classification of [Anselmi:2021rye] to $f(R)$ theories directly in the Jordan frame, which allows us to identify the correct definition for $\alpha$. In particular, assuming that the beta function be a power series, we classify the models and derive the leading predictions class by class without specifying the form of $f(R)$. Moreover, starting from a beta function with generic coefficients, we work out the expression of all the $f(R)$ that belong to the same class in terms of such coefficients.
@@
Comparing the results with the most up-to-date experimental constraints we find conditions on the first nonvanishing coefficient of the beta functions and exclude all classes but two. Remarkably, if we focus on polynomial $f(R)$, which in general relativity are generated by renormalization anyway, we find that the only one suitable for slow-roll inflation is the Starobinsky model. In fact, extending the analysis to theories with a linear beta function, it turns out that all the other polynomials belong to such a class. A particular property of these theories is that they cannot realize slow-roll inflation but only constant-roll inflation [Motohashi:2014ppa], i.e. an accelerated expansion with a constant rate. We highlight some difficulties in interpolating between slow-roll and constant-roll inflation, and, thus, between the Starobinsky model and its polynomial extensions.
@@
Finally, once we derive all the quantities in the Jordan frame we provide the map to the Einstein frame for comparison with [Anselmi:2021rye].
@@
The paper is organized as follows. In section 2 we review $f(R)$ theories, introduce the RG analogy and find the definition of $\alpha$ that simplifies the equations. In section 3 we derive the quadratic action for tensor and scalar perturbations in a generic $f(R)$ theory and recall the derivation of the spectra used in [Anselmi:2021rye]. In section 4 we work out the leading-order predictions for the spectral indices and the tensor-to-scalar ratio and show that every class that does not have quadratic or cubic beta function is ruled out by data. In section 5 we derive the general form of the $f(R)$ that belongs to the surviving classes and their power spectra. In section 6 we derive the perturbative map between the Jordan frame and the Einstein frame. In section 7 we study a special class of theories with linear beta function and point out the issues with these models. Finally, section 8 contains our conclusions.
@@
Notation and conventions: We use the signature (+,,,)(+,-,-,-) for the metric tensor. The Riemann and Ricci tensors are defined as $R^{\mu}_{\ \nu\rho\sigma}=\partial_{\rho}\Gamma^{\mu}_{\nu\sigma}-\partial_{\sigma}\Gamma^{\mu}_{\nu\rho}+\Gamma^{\mu}_{\alpha\rho}\Gamma^{\alpha}_{\nu\sigma}-\Gamma^{\mu}_{\alpha\sigma}\Gamma^{\alpha}_{\nu\rho}$ and $R_{\mu\nu}=R^{\rho}_{\ \mu\rho\nu}$, respectively. In the case of cosmological perturbations we omit the integral over the space momentum and use the word “action” for the integral over time of the lagrangian density.
@@ 2 $f(R)$ theories and the cosmic RG flow
We start by recalling the action and equations of motion on the Friedmann-Lemaître-Robertson-Walker (FLRW) background. The starting action reads
@@
where $G$ is the Newton constant. The equations of motion are
@@
where $H\equiv\dot{a}/a$ is the Hubble constant. Now we adapt the procedure of [Anselmi:2021rye] to $f(R)$ theories. The ingredients to define a RG flow are as follows:
@@
    a first-order differential equation for that small quantity (analogous of the beta function);
@@
Indeed, this is a good choice if we work with a specific $f(R)$ model (see [Anselmi:2020lpp]). However, we would like to treat these types of theories on general grounds. It turns out that another useful quantity is instead
@@
which we show below to be small during inflation. Following [Anselmi:2021rye] we also define
@@
$v\beta_{\alpha}=\alpha(1+2\alpha)-(1-\alpha^{2})\varepsilon,$ (2.10)
@@
where the $b_{i}$ are constant coefficients. Note that if we admit a linear beta function, then $\varepsilon=\mathcal{O}(\alpha)$, while $\eta_{H}=\mathcal{O}(\alpha^{0})$. In this case slow-roll inflation cannot be realized because $\eta_{H}$ is constant in the far past. This class of models is discussed in section 7.
As mentioned above, we could have chosen $\varepsilon$ as a small quantity to define the RG flow. In that case we would have had
@@
Again, assuming $\varepsilon\ll 1$, we need that $\beta_{\varepsilon}$ be at least quadratic to guarantee $\eta_{H}=\mathcal{O}(\varepsilon)$. Since $\varepsilon(\alpha)$ is a simple power series, it means that also $\alpha(\varepsilon)$ is. We conclude that assuming $\alpha\ll 1$ and $\beta_{\alpha}=\mathcal{O}(\alpha^{2})$ is equivalent to assume $\varepsilon\ll 1$ and $\beta_{\varepsilon}=\mathcal{O}(\varepsilon^{2})$. This means that, during the inflationary phase, the roles of $\alpha$ and $\varepsilon$ are interchangeable without loss of generality. Passing from one to the other is just a perturbative redefinition.
Finally, the background spacetime needs to reach the de Sitter one in the asymptotic past. This means that the first nonvanishing coefficient of the beta function must be negative. Because of these properties, the whole construction resembles that of beta functions and running-coupling constants in quantum field theory, where an asymptotically-free theory requires a negative beta function. Here the “flow” is represented by the background evolution during the perturbative phase of inflation. More details can be found in [Anselmi:2021rye]. It is important to highlight that such an analogy is just a mathematical correspondence and has nothing to do with quantum corrections, which are not considered in this paper.
@@
Equation (2.17) can be solved by a simple power series
@@
Equation (2.25) can be solved by truncating the beta function. Since its formal expressions do not depend on the frame, we can use the same listed in [Anselmi:2021rye]. For class I and II at the leading-log order (LLO) we have
@@
We do not report the next-to-leading-log (NLL) corrections since in this paper we can directly use the results of [Anselmi:2021rye]. However, it is important to highlight that all the results for power spectra and spectral indices will eventually be written as expansions in $\alpha_{k}$, which is fixed at a pivot scale $k_{*}$ by the experimental data. Then thanks to the second equation in (2.26) we can evolve $\alpha$ to an arbitrary $k$. For example, in class I we have
@@
Since $\mathcal{P}$ can be viewed as a function of $\alpha(\ell)$ and $\alpha_{k}$, its conservation on superhorizon scales means that the dependence on $\alpha$ drops and $k$ appears only through $\alpha_{k}$. More detail can be found in [Anselmi:2020shx].
@@ 3 Perturbations
In this section we outline the general strategy to derive the quadratic action for the perturbations and solve their equations of motion by briefly reviewing the method used in [Anselmi:2021rye]. However, we can directly export the solutions of the Mukhanov-Sasaki equation as well as the expressions for the power spectra and spectra indices at the NNLL and NLL for tensor and scalar perturbations, respectively. The reason is that in [Anselmi:2021rye] those results were derived for a generic action quadratic in the perturbations, without specifying its origin.
@@
where $\sigma$ is a function that depends on background quantities. Typically [Anselmi:2021rye], $\sigma$ is analytic in $b_{i}$, which means that it can only depend on $v$ and $\beta_{\alpha}$, as well as on $\alpha$ explicitly, because $H$, $F$, $f$ and $R$ are different for each class and not analytic in the first nonvanishing $b_{i}$.
@@
where $g_{n}$ are functions that depend on $\sigma$ and on the $w_{m}$ with $m<n$. The Bunch-Davies condition
@@
Finally, the solution in terms of $f_{n}$ is [Anselmi:2020shx]
@@
The procedure just explained gives the $w$ as an explicit function of $\eta$. However, it is possible to single out the part of $w$ (more precisely of $\eta w$) that survives the superhorizon limit. Such part depends on $\ln\eta$ only. For this reason, it can be written as a function of $\alpha$. We summarize the steps here below. More details can be found in [Anselmi:2020shx] and  [Anselmi:2021rye].
@@
and equation (3.9) turns into
@@
Both in the tensor and scalar case, the constant $C$ can be chosen such that $w^{\text{C}}$ coincides with the usual gauge-invariant perturbations, which are also conserved in the superhorizon limit. The power of this method is that it gives a recipe to obtain gauge-invariant perturbations that are conserved on superhorizon scales by construction. Although this result might not seems surprising in the case of single-field inflation, it becomes particularly relevant in the case of multifield inflation, where it is nontrivial to find conserved perturbations. See [Anselmi:2021dag] and [Grzadkowski:2025yxr] for applications of the RG techniques in the context of double-field inflation.
@@
As anticipated in the beginning of this section, since in [Anselmi:2021rye] the functions $w_{n}$, as well as $J$ and $Q_{0}$, have been computed assuming a power series for $\sigma$ with generic coefficients, we can directly use those results. The expressions for $w_{n}$ are analytic in $b_{i}$, while those for $J$, and therefore those for power spectra, depend on the class. The reason is that the transformation from $w$ back to the original perturbation involves $F$, $H$ and $f$.